OpenStudy (anonymous):

A boardwalk that is x feet wide is built around a rectangular pond. The combined area of the pond and the boardwalk is 4x^2 + 140x + 1200 square feet. What are the dimensions of the pond?

OpenStudy (anonymous):

4x^2 + 140x +1200 factor it out first: 2 *2(x^2 + 35x+ 300) = 4(x+15)(x+20) =( 2x+30) (2x+40) After simplifying area of the pond and sidewalk equates to: (2x+30)(2x+40) You know that x=sidewalk, if you want to exclude sidewalk set x = 0, to find area of only the pond = (2(0) (30)) (2(0) + 40) = 30 (40) remember Area of rectangle = base times height? Area without sidewalk = 30 times 40 so the dimensions is 30 by 40